Which binomial is a factor of 3X^2+2x-5
Answer Choices:
A 3x-1
B x-1
C 3x-5
D x-5

Answer by checkley71(8403)

3x^2+2x-5=0
(3x+5)(x-1)=0
ANSWER B) (X-1)